OpenStudy (anonymous):

Which shows 602 - 402 being evaluated using the difference of squares method? a. 602 - 402 = (3600 + 1600)(3600 - 1600) = 10,400,000 b. 602 - 402 = 3600 - 1600 = 2,000 c. 602 - 402 = (60 - 40)2 = 202 = 400 d. 602 - 402 = (60 + 40)(60 - 40) = (100)(20) = 2,000 medals given!

OpenStudy (compphysgeek):

assuming that 602 - 402 means \(60^2 - 40^2 \) answer d is correct. \[ (a + b) \times (a - b) = a^2 - b^2\]